8-e=2e-16;8 is this a solution of the equation
Answer by rapaljer(4671)

Check it out and see if it is! Substitute e=8 into the equation and see if it is true.

8-e = 2e - 16
8 - 8 = 2(8) - 16
0=0

True, so it is a solution!

R^2 at SCC